Calculate the vapor pressure of 3-pentanone (b.p. 102.0 °C) at the gas chromatography column temperature of 96.0C using the form of the Clausius-Clapeyron equation shown below, P1 P. vap where R is the ideal gas constant, ???? is the enthalpy of vaporization, T1 and T2 are two different temperatures, and P1 and P2 are the vapor pressures at the respective temperatures. The enthalpy of vaporization can be estimated using Troutons rule, ??vap-(88 J mol-1 K-1), Tp. Number Torr Calculate the vapor pressure of pentane (b.p. 36.0 °C) at the same column temperature. Number Torr For compounds of the same chemical class, what is the relationship between vapor pressure of a compound and retention in gas chromatography?
Answer #1

1.

ln(P2 / 760) = ( 88 / 8,314 ) * [ ( 1 / 375.15 ) - ( 1/ 369.15 ) ]

P2 = 759.65 torr

2.

ln(P2 / 760) = (88 / 8.314) * [ ( 1/ 309.15 ) - ( 1 / 369.15 ) ]

P2 = 764.24 torr

So, the relation ship is

Vapour pressure \alpha ( 1 / retention time )

SO, (c) As the vapour pressure decrease , retention time increases.